OpenStudy (anonymous):

Fan and medal for easy problem check below :) Solve the following system of equations. -3x + 4y = 80 x + 4y = 32 A. x = 2, y = 21.5 B. x = 11, y = -12 C. x = 8, y = 6 D. x = -12, y = 11

OpenStudy (jadedinsomniac):

i got this hold up

OpenStudy (anonymous):

\[-3x+4y=80, x+4y=32\] multiplying the second equation by (-1) we get \[-3x+4y=80, -x-4y=-32\] now, we sum both equations getting \[-4x=48 \rightarrow x=-12\] and replacing x=-12 in the first equation we get \[-3(-12)+4y=80 \rightarrow 36+4y=80 \rightarrow 4y=44 \rightarrow y=11\] so x=-12 and y=11 is the solution to the system of equations